If 15 pounds of fertillizer will cover 1500 square feet of lawn, how much fertilizer is needed to cover 2500 square feet?
It would require ______ pounds to cover 2500 square feet.

Thanx

Answer by fractalier(2101)

Set up a proportion, cross-multiply and solve for x...here goes...
15 / 1500 = x / 2500 (you should be able to see the answer 25 already, but...)
1500x = 37500
x = 37500/1500
x = 25 pounds